OpenStudy (anonymous):

I'm in CA, Geometry A. I need the answers to Unit 3, Lesson 10 Parallel and Perpendicular lines test. Only have 4 days to finish everything & i'm behind so if anyone has the answers or could at least help I need it ASAP!

jigglypuff314 (jigglypuff314):

Please take some time to read this site's code of conduct http://openstudy.com/code-of-conduct Open Study has a policy for "Give Help, Not Answers". Violations to Open Study's policies could result in a possible account suspension. Therefore it would be best if you wouldn't request answers like so. I'll do my best to help you though, if I can :)

OpenStudy (anonymous):

Oh ok i'm sorry! 1. Which diagram suggests a correct construction of a line parallel to given line 1 and passing through given point Q?

jigglypuff314 (jigglypuff314):

http://www.mathopenref.com/constparallel.html

OpenStudy (anonymous):

I'm so confused :/

jigglypuff314 (jigglypuff314):

if you're using angle copy method, it would look more like option B where there are two angles along the line going from line 1 to point Q

OpenStudy (anonymous):

Okay thank you i'll try to figure it out. I'm not good at this kinda stuff .. 2. Find the value of x and y. A. x=15,y=12 B. x=14,y=11 C. x=14,y=12 D. x=15,y=11

OpenStudy (ranga):

4x+2 and 62 are alternate interior angles and therefore they are equal. Solve for x. 12y and 144 are alternate interior angles and therefore they are equal. Solve for y.

jigglypuff314 (jigglypuff314):

opposite interior angles are congruent (equal to each other) so 4x + 2 = 62 12y = 144

OpenStudy (anonymous):

I'm not even sure how to solve for x and y honestly.... i'm so dumb.

jigglypuff314 (jigglypuff314):

don't be so pessimistic, I'm sure you'll be able to understand and learn how :) for 4x + 2 = 62 subtract both sides by 2 so -2 -2 and you get 4x = 60 then divide both sides by 4 /4 /4 x = 60/4 = ___ then for y you would divide both sides by 12 so y = 144/12 = __

OpenStudy (anonymous):

x=15 and x=12, is that right?

OpenStudy (anonymous):

***x=15, and y=12

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

yayyy . 3. What is the relationship between angle 8 and angle 4? A. corresponding angles B. alternate exterior angles C. same-side interior angles D. alternate interior angles

jigglypuff314 (jigglypuff314):

by definition, they would be corresponding angles http://www.mathsisfun.com/geometry/corresponding-angles.html

OpenStudy (anonymous):

4. Which construction is illustrated above? A. a perpendicular to a given line from a point on the line B. a line segment congruent to a given line segment C. the perpendicular to a given line from a point on the line D. the perpendicular bisector of a line segment

jigglypuff314 (jigglypuff314):

Not B and D because there are no line segments involved... the difference between A and C is "a perpendicular" and "the perpendicular" (I'm not sure if this is right...but) given one line and one point, there would be only one perpendicular. therefore C

OpenStudy (anonymous):

Thank you! My computer is about to die so I gotta get off :/ Can we continue later?

jigglypuff314 (jigglypuff314):

if I can :) have a nice day :)

OpenStudy (anonymous):

You on?

OpenStudy (anonymous):

I finished this myself! Got 9/12 for the multiple choice questions and waiting on the teacher to grade the type in questions. Thank you for helping :) Gonna close this question.